A block of ice with mass 2.00 kg slides 0.750 m down an inclined plane that slopes downward at an angle of 36.9 degrees below the horizontal. If the block of ice starts from rest, what is its final speed? You can ignore friction.
Answer #1
Concepts and reason

The concepts required to solve this problem are Newton’s second law and kinematics equation of motion.

Initially, Draw the free body diagram from given data. Then find the acceleration of block of ice on inclined plane by using Newton’s second law of motion. Finally, calculate the final speed of block by using kinematic equation.

Fundamentals

The Newton’s second law states that the net force on an object is the product of mass of the object and final acceleration of the object. The expression of newton’s second law is,

F=ma\sum {F = ma}

Here, F\sum F is the sum of all the forces on the object, mm is mass of the object, and aa is the acceleration of the object.

The kinematics equation of motion is,

vf2=vi2+2adv_{\rm{f}}^2 = v_{\rm{i}}^2 + 2ad

Here, dd is the distance travelled by the object, vi{v_{\rm{i}}} is the initial velocity, vf{v_{\rm{f}}} is the final velocity, and aa is the acceleration.

Draw the following free body diagram of block from given data:

mg cos e
mg sine
mg

Here, m is mass of block, g is acceleration due to gravity, N is normal force on the block, and θ\theta is angle made by incline with horizontal.

The acceleration a of the block on the incline plane is equal to gsinθ.g\sin \theta .

a=gsinθa = g\sin \theta

Use the following equation of kinematics to find the final speed of block.

vf2=vi2+2adv_{\rm{f}}^2 = v_{\rm{i}}^2 + 2ad

Here, vf{v_{\rm{f}}} is final speed, vi{v_{\rm{i}}} is initial speed, a is acceleration, and d is distance.

Substitute gsinθg\sin \theta for aa in the above equation and solve for final speed vf{v_{\rm{f}}} .

vf2=vi2+2(gsinθ)dvf=vi2+2(gsinθ)d\begin{array}{c}\\v_{\rm{f}}^2 = v_{\rm{i}}^2 + 2\left( {g\sin \theta } \right)d\\\\{v_{\rm{f}}} = \sqrt {v_{\rm{i}}^2 + 2\left( {g\sin \theta } \right)d} \\\end{array}

Substitute 0m/s0{\rm{ m/s}} for vi{v_{\rm{i}}} , 9.8m/s29.8{\rm{ m/}}{{\rm{s}}^2} for gg , 36.936.9^\circ for θ\theta , and 0.750m0.750{\rm{ m}} for dd in the equation vf=vi2+2(gsinθ)d{v_{\rm{f}}} = \sqrt {v_{\rm{i}}^2 + 2\left( {g\sin \theta } \right)d} and calculate vf{v_{\rm{f}}} .

vf=(0m/s)2+2((9.8m/s2)sin36.9)(0.750m)=2.97m/s\begin{array}{c}\\{v_{\rm{f}}} = \sqrt {{{\left( {0{\rm{ m/s}}} \right)}^2} + 2\left( {\left( {9.8{\rm{ m/}}{{\rm{s}}^2}} \right)\sin 36.9^\circ } \right)\left( {0.750{\rm{ m}}} \right)} \\\\ = 2.97{\rm{ m/s}}\\\end{array}

Ans:

The final speed of the block of ice is 2.97m/s2.97{\rm{ m/s}} .